The solubility of CaCO_3 is pH dependent, Calculat
Answer #1

CO32- (aq) + H2O \rightleftharpoons HCO3- + OH- Carbonate ion hydrolyses in water to form a basic solution.

Kb =[HCO3-] [OH-]/[CO3-] Therefore, [HCO3-][OH-] = Kb[CO3-]

Kb(CO32-) = 2.1 x10-4

For this equation, CaCO3(s)+H2O(l)⇌Ca2+(aq)+HCO3−(aq)+OH−(aq)

equilibrium constant K = [Ca2+][HCO3-][OH-] CaCO3 is in the solid state, and water is in excess. Hence their concentration is not considered in the equilibrium expression

K = [Ca2+][HCO3-][OH-] =    K = [Ca2+] Kb[CO3-] = [Ca2+][CO3-] x2.1x10-4  = Kspx2.1x10-4

= 4.5x10-9x 2.1x10-4 = 10.29x10-13 = 1.029x10-12  = 1.0x10-12

D. CaCO3(s) → ← Ca2+ + CO3 2-

Ksp = [Ca2+] [CO3 2-] (1) If you don't take into account the hydrolysis of carbonate ion

S = [Ca2+] = [CO3 2-] = Ksp = 6.71x10-5 (Has been given correct)

Taking into account Ka2 K2 = [H+ ] [CO3 2-]/ [HCO]3-

S = [Ca2+] = [CO3 2-] + [HCO3 -] = [CO3 2-] + [H+ ] [CO3 2-]/ K2

S = [Ca2+] = [CO3 2-] ( 1 + [H+ ] /K2 )

Multiply both sides by Ca+2 ions

S2= [Ca2+] 2 = Ksp ( 1 + [H+ ] /K2) Substituting the values of Ksp and K2 = 5.0x10-11 and[ H+]

We can get the solubility of CaCO3 at any Desired pH

For instance at 8.3 pH

-log[H+] = 8.3 [H+] = 5.01x10-9

S2 = 45.45x10-8   S = 6.74x10-4 mol L-1